OpenStudy (anonymous):

find equation its graph is a parabola with x-intercepts 2 and -1 and y-intercept 6

OpenStudy (anonymous):

right

OpenStudy (anonymous):

Equation will have the form: \[y=ax^2+bx+c\]With the information given. y intercept is 6 means that y=6 when x=0----->c=6 x-intercept =2 and -1 are the solutions when y=0. Then: \[a(2)^2+b(2)+6=0 \rightarrow 4a+2b=-6\]and\[a(-1)^2+b(-1)+6=0 \rightarrow a-b=-6\] Solving the system: 4a+2b=-6 a-b=-6 will give you the a and b you need to complete the function